OpenStudy (anonymous):

Put the equation x2+y2+ax+by+c=0 in standard form.

OpenStudy (anonymous):

Put the equation x2+y2+ax+by+c=0 in standard form.

OpenStudy (ybarrap):

m^2+2 m x+n^2+2 n y+x^2+y^2 = p^2 where (x+m)^2+(y+n)^2=p^2 In your case, a=2m b=2n c= m^2 + n^2 - p^2 so p^2 = m^2 + n^2-c)=(a/2)^2 + (b/2)^2 - c So x2+y2+ax+by+c =(x+a/2)^2+(y+b/2)^2=(a/2)^2+(b/2)^2 -c, which is in standard form Where the radius = p and the center of the circle is at point: <-m,-n>.

OpenStudy (anonymous):

I don't understand